In a university one-half (50%) of the entering freshman attend two semesters a year, pay $12,000. The university earns a contribution to overhead on revenues of 30%. (To simplify, assume that no students are allowed to transfer and all enrolled students enter as freshman.)

What would be the value to the university's contribution of a retention program that increased the 5-year retention rate from 50% to 60%?
Answer #1

Fee (P) = $ 12,000

Contribution margin (CM) = 30 %

Frequency (RF) = 5

Average defection rate (DR) = 1 - 50% = 50 %

Buyer's Life cycle (BLC) = 1/DR = 1/50% = 2

Value of a Loyal Customer (VLC) = P*CM*RF*BLC = 12000*30%*5*2 = $ 36,000

After retention rate increase to 60%, DR = 1-60% = 40%

BLC = 1/40% = 2.5

VLC = 12000*30%*5*2.5 = $ 45,000

Value to the university of retention program = 45000 - 36000 = $ 9,000